OpenStudy (anonymous):

differentiate 5^x/(ln5)

OpenStudy (anonymous):

\[\frac{ 5^x }{ \ln5 }\]

OpenStudy (anonymous):

Quotient Rule: d/dx(f(x)/g(x)=gf'-fg'/g^2 So, d/dx= 5^\[(5^xlog(5)\log(x)-(1/x)5^x)/\log(x)^2\]

OpenStudy (loser66):

check ,please. It's not good to me.

OpenStudy (loser66):

we don't apply quotient rule since ln5 is a constant.

OpenStudy (anonymous):

okay thanks so much :)

OpenStudy (anonymous):

my fault thought said ln(x)

OpenStudy (anonymous):

its just 5^x

OpenStudy (anonymous):

okay, so then how do you solve it without the quotient rule?

OpenStudy (anonymous):

d/dx (5^x)=ln(5)d^x ln(5)5^x/ln(5) is 5^x

OpenStudy (loser66):

I just retype Standish's work \[\left(\frac{5^x}{ln5}\right)'= \frac{1}{ln5} (5^x)'= \frac{1}{ln5} ln5 *5^x= 5^x\]

OpenStudy (anonymous):

gotcha thanks :)